OpenStudy (anonymous):

Solve for x: x+2 = 24/x

OpenStudy (anonymous):

okay, let's cross multiply first and you'll get : x(x+2)= 24 so now you can expand it and get : \[x^2+2x-24 =0\] \[(x-4)(x+6) = 0\] so, x = 4 and x = -6 ^_^

OpenStudy (anonymous):

AHH youare so helpful :) so four down from this prob. is one i am stuck on as well.

OpenStudy (anonymous):

glad I could help ^_^

OpenStudy (anonymous):

the problem four down from this one i titled it simplify with steps [please]

OpenStudy (anonymous):

PLEASE :)

OpenStudy (anonymous):

LOL alright!